Ministry of Education put a guideline that maximum of 30% students 

are allowed to be in campus to control the spread of Covid-19. In average, 13 students 

will come to Statistics class with standard deviation of 3 students. A sample of 20 

students are chosen. Find probability that the average number of students come to 

Statistics class is at least 18?


1
Expert's answer
2021-05-10T13:38:20-0400

"\mu = 13 \\

\sigma = 3 \\

n = 20 \\

P(\bar{x}>18) = 1 -P(Z<\frac{\bar{x}- \mu}{\sigma / \sqrt{n}}) \\

=1 -P(Z<\frac{18-13}{0.6711}) \\

= 1 -P(Z<7.45) \\

= 1 -0.9999 \\

= 0.001"